_     __
7Ö3 - 5Ö75

Answer by AnlytcPhil(1806)


  _     __
7Ö3 - 5Ö75

Break 75 into its prime factors: 
75 = 3·25 = 3·5·5 = 3·5²,
so replace 75 by 3·5²
  _     ____
7V3 - 5Ö3·5²

Take individual square roots of 3 and of 5²:
  _     _  __
7Ö3 - 5ÖÖ5²

The square root of 5² is just 5
  _     _ 
7Ö3 - 5Ö3·(5)
  _      _ 
7Ö3 - 25Ö3
            _
Factor out Ö3
 _       
Ö3(7 - 25)
  
Combine what's in the parentheses
 _       
Ö3(-18)
    _
-18Ö3

Edwin
AnlytcPhil@aol.com